# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v4

@@ Abstract
We derive formulae that resum, at a given order in the soft limit, the infinite series of Post-Minkowskian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in the Schwarzschild geometry in the probe limit. Specifying to the case of circular orbits, we compute the waveform and the energy flux to order 30PN, and compare it against the available results in the literature. Our results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a simple mathematical proof of the Heun connection formula.
@@ 1 Introduction
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms, in the probe limit, were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
@@
More recently, multipole methods Bini:2023fiz ; Bini:2024ijq ; Bini:2024rsy , scattering amplitude techniques Jakobsen:2021smu ; Mougiakakos:2021ckm ; Cristofoli:2021vyo ; DiVecchia:2023frv ; Brandhuber:2023hhy ; Herderschee:2023fxh ; Elkhidir:2023dco ; Georgoudis:2023lgf ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Georgoudis:2023eke ; Georgoudis:2023ozp ; Georgoudis:2024pdz ; Brunello:2024ibk ; Alessio:2024onn and black hole perturbation theory Fucito:2024wlg have been applied to the study of gravitational scattering processes, fostering a fruitful interplay and mutual cross-checking among these complementary approaches. Within the quantum field theory framework, the PM series describes the expansion in loops (weak coupling) of the scattering amplitude keeping finite the relative velocity of the binary. Using perturbation theory, we will show that at any fixed order in the gravitational coupling, the entire series of PN corrections in the probe limit of the gravitational waveform can be resummed and expressed in terms of a single hypergeometric function. Analogous results hold for the PN expansion itself, where the exact dependence on the gravitational coupling corrections is determined at any given order in the small-velocity limit.
@@
The PM and PN series provide novel hypergeometric representations of the confluent Heun function along the lines of the well studied MST solutions Mano:1996vt . The main difference between the two, is that the MST ansätze involves an infinite sum of hypergeometric functions, while the solution we put forward here is written in terms of a single hypergeometric function and its first derivative. The crucial advantage of this representation is that it provides a simple mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The connection matrix coefficients are given as ratios of gamma functions with arguments depending on a single (but highly non-trivial) function codifying the holonomies of the confluent Heun function around the singular point. This function, known as the "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk , can be related to the so called Seiberg-Witten period of a ${\cal N}=2$ supersymmetric gauge theory Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe or the "anomalous dimension" in the effective field theory description of the gravitational wave Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions (and its confluences) from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resumming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe the PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to 30 PN orders. This result will be tested against the computation at 22 PN orders available in literature Fujita:2012cm and the numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST approach. The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. Section 5 contains our conclusions. In Appendix A we derive the PM and PN expansions of the non-confluent Heun function.
@@
The second order ordinary differential equation called the confluent Heun equation is introduced in (6). It has two regular singularities at $z=1,\infty$ and an irregular one at $z=0$. The equation is specified by five complex parameters dubbed here as $\{m_{1},m_{2},m_{3},u,x\}$. The expansion of the general solution of this equation in the ”soft limit” $x\ll 1$ is derived in Section 2. These results apply to any physical problem governed by an equation of confluent Heun type. The focus of this paper is on the study of gravitational wave forms in the Schwarzschild geometry. The Heun-gravity dictionary in this case reads
@@ 2 The confluent Heun equation
${\rm HeunC}^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right){\rm HeunC}^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}{\rm HeunC}(t)=0$ (2)
@@
${\rm HeunC}[\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},\lambda,t]$ (3)
$t^{1-\beta_{0}}\text{HeunC}[\alpha_{0}+(1-\beta_{0})(\lambda-\beta_{1}),\alpha_{1}+\lambda(1-\beta_{0}),2-\beta_{0},\beta_{1},\lambda,t]$
@@ 2.1 Confluent Heun-Gauge dictionary
with $\delta=m_{1}m_{2}{+}m_{1}m_{3}{+}m_{2}m_{3}{-}m_{1}{-}m_{2}{-}m_{3}$, $x$ the gauge coupling and $u$ the Coulomb branch parameter. (2) and (6) are related by the Heun-gauge dictionary
$G_{1}(z)$ $=$ $z^{-{\alpha_{1}\over\lambda}}{\rm HeunC}({\textstyle\frac{1}{z}})\qquad,\qquad t=z^{-1}\qquad,\qquad\lambda=-x$
@@
Equivalently, the confluent Heun equation can be written as a differential equation in the variable $y=x/z$, taking the form
@@
$g(x)G_{0}(y)=G_{1}({\textstyle\frac{x}{y}})=({\textstyle\frac{y}{x}})^{{\alpha_{1}\over\lambda}}{\rm HeunC}({\textstyle\frac{y}{x}})$ (9)
and $g(x)$ an arbitrary constant. We will use both descriptions (6) and (8) of the confluent Heun equation, and we will refer to them as CHE. In particular, we will focus on the soft (or large wavelength) limit where $x$ is taken small keeping finite either $z$ or $y$. In the gravity context, the two limits correspond to the PM and PN expansions of the confluent Heun solution, so we will refer to $G_{0}$ and $G_{1}$ as the PM and PN representations of the confluent Heun function respectively. In particular, sending $x\to 0$, (8) and (6) reduce to hypergeometric equations with solutions
@@
with $\alpha=\pm$ labelling the two independent solutions,
@@
and $a\underset{x\to 0}{\approx}\sqrt{u}$. On the other hand, at order $k$ in $x$,  [^3] This is the only expansion that will appear in our formulae. With an abuse of notation we will define our results a PN expansion at order $k$ even if other contributions to our final formulae will be given in closed form. we will look for solutions of the form
${\rm PM}_{k}:$ $\qquad G^{0}_{\alpha}(y)=y^{\frac{1}{2}-\alpha a-m_{3}}\sum_{i=0}^{k}\sum_{j=0}^{\infty}c^{PM}_{ij}(\alpha a)x^{i}y^{j-i}$
${\rm PN}_{k}:$ $\qquad G^{1}_{\alpha}(z)=z^{-\frac{1}{2}+\alpha a+m_{3}}\sum_{j=0}^{k}\sum_{i=0}^{\infty}c^{PN}_{ij}(\alpha a)z^{i-j}x^{j}$ (12)
where $a(u)$ is a function of $u$. Writing
@@
one can determine the coefficients $c^{PM}_{ij}(a)$, $c^{PN}_{ij}(a)$ and $u_{i}(a)$ by solving the CHE equations order by order in $x$. We notice that $a$ codifies the monodromy of the solutions around the irregular point $z=0$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.
@@ 2.2 PM approximation: exterior region
It is easy to check that $c_{ij}$, $\widehat{c}_{ij}$ are invariant under $a\to-a$. This is true at any PM order, so all the dependence on $\alpha$ of the solution is codified in $H^{0}_{\alpha}$. The results (19) are special cases of more general relations that determine the coefficients $c_{i,i-1}$, $\widehat{c}_{i,i-1}$ in terms of $u$
@@
$\gamma_{m}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{\sum_{i=1}^{\infty}{\partial_{m_{3}}u_{i}\over i}x^{i}\right\}\quad$ (20)
The combination in the exponential will be denoted in the following as
${\cal F}_{\rm inst}(a,x)\equiv-\sum_{i=1}^{\infty}{u_{i}(a)\over i}x^{i}$ (21)

### v5

@@ Abstract
We derive formulae that resum, at a given order in the soft limit, the infinite series of Post-Minkowkian (small gravitational coupling) or Post-Newtonian (small velocities) corrections to the gravitational waveform produced by particles moving along a general (open or closed) trajectory in Schwarzschild geometry in the probe limit. Specifying to the case of circular orbits, we compute the wave form and the energy flux to order 30PN, and compare it against the available results in the literature. The results are based on a novel hypergeometric representation of the solutions of the Heun equation (and its confluence), that leads to a simple mathematical proof of the Heun connection formula.
@@ 1 Introduction
In the case of circular orbits, the PM and PN approximations are linked and characterised by a single expansion parameter $v$ (the tangential velocity). The current state-of-the-art for the emitted energy in perturbation theory is the 22nd PN (order $v^{44}$) computation in Fujita:2012cm . The expansion is known to be converging slowly Cutler:1993vq ; Poisson:1993vp due to the presence of logarithmic terms, leading to an infinite tower of tails and tails of tails non-local interactions Damour:2008gu ; Damour:2009kr ; Buonanno:2009zt ; Damour:2012ky . In Aminov:2020yma ; Bianchi:2021xpr ; Bianchi:2021mft ; Bonelli:2021uvf ; Bonelli:2022ten ; Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe novel techniques, capitalizing on recent advances in quantum field theories, have been exploited to derive an exact formula for the Heun connection matrix and alternative formulations of the waveforms were proposed in terms of the instanton partition function of a SU(2)2 quiver gauge theory. Following this path the infinite tower of logarithmic terms have been resummed into exponentials.
@@
More recently, multipole methods Bini:2023fiz ; Bini:2024ijq ; Bini:2024rsy , scattering amplitude techniques DiVecchia:2023frv ; Brandhuber:2023hhy ; Herderschee:2023fxh ; Elkhidir:2023dco ; Georgoudis:2023lgf ; DeAngelis:2023lvf ; Brandhuber:2023hhl ; Georgoudis:2023eke ; Georgoudis:2023ozp ; Georgoudis:2024pdz ; Brunello:2024ibk ; Alessio:2024onn and black hole perturbation theory Fucito:2024wlg have been applied to the study of gravitational scattering processes, fostering a fruitful interplay and mutual cross-checking among these complementary approaches. Within the quantum field theory framework, the PM series describes the expansion in loops (weak coupling) of the scattering amplitude keeping finite the relative velocity of the binary. Using perturbation theory, we will show that at any fixed order in the gravitational coupling, the entire series of PN corrections to the probe limit of the gravitational waveform can be resummed and expressed in terms of a single hypergeometric function. Analogous results hold for the PN expansion itself, where the exact dependence on the gravitational coupling corrections is determined at any given order in the small-velocity limit.
@@
The PM and PN series provide novel hypergeometric representations of confluent Heun function closely in spirit to the well studied incoming and upgoing MST solutions Mano:1996vt . The crucial difference between the two, is that while the MST ansatz involves an infinite sum of hypergeometric functions, our solution is written in terms of a single hypergeometric function and its first derivative. The crucial advantage of this representation is that it provides a simple mathematical proof of the Heun connection formulae originally found in Aminov:2020yma ; Bianchi:2021xpr ; Bonelli:2021uvf ; Bonelli:2022ten . The connection matrix coefficients are given as ratio of gamma functions with arguments depending on a single (but highly non-trivial) function codifying the holonomies of the confluent Heun function around the singular point. This function, known as the "renormalized angular momentum" in the MST context Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk , can be related to the so called Seiberg-Witten period of a ${\cal N}=2$ supersymmetric gauge theory Consoli:2022eey ; Bautista:2023sdf ; Aminov:2023jve ; ToVSapQNM ; Fucito:2023afe or the "anomalous dimension" in the effective field theory description of the gravitational wave Goldberger:2009qd ; Trestini:2023wwg ; Almeida:2021jyt ; Edison:2023qvg . The Heun connection formula has been recently applied in a variety of physical contexts, including the study of QNMs for gravity backgrounds (BHs, fuzzballs, D-branes), correlators in thermal field theories, holography, cosmological perturbations and scattering amplitudes in gravity Bianchi:2021xpr ; Bianchi:2021mft ; Fioravanti:2021dce ; Bianchi:2022wku ; Dodelson:2022eiz ; Dodelson:2022yvn ; Imaizumi:2022qbi ; Fioravanti:2022bqf ; Imaizumi:2022dgj ; Bianchi:2022qph ; Dodelson:2023vrw ; Bianchi:2023rlt ; Bianchi:2023sfs ; Giusto:2023awo ; Aminov:2023jve ; BarraganAmado:2023apy ; Lei:2023mqx ; BarraganAmado:2023wxt ; Fucito:2023afe ; Bautista:2023sdf ; Ge:2024jdx ; Cipriani:2024ygw ; Arnaudo:2024rhv ; Bianchi:2024vmi ; Bianchi:2024mlq ; DiRusso:2024hmd ; Bianchi:2024rod ; Aminov:2024aan ; BarraganAmado:2024tfu ; Bautista:2024emt ; Liu:2024eut ; Silva:2025khf ; Aprile:2025hlt ; Bhatta:2025kil .
@@
To keep the material self-contained, in this paper we present a derivation of the Heun solutions (and its confluences) from first principles with no reference to supersymmetric gauge theory or localization, which sit in the background as inspiration to some of our ansätze, that nonetheless could be also accepted as such. Finally, we remark, that even if the main motivation for resuming PM and PN series relies on the study of open orbits, our hypergeometric representations provide an algorithm which is very efficient when implemented in a Mathematica code to describe PN expansions of bounded orbits. Indeed, here we compute the gravitational waveform and the energy flux, in the case of circular orbits, up to the 30th PN order [^2] The ancillary mathematica files coming with this paper should help the reader to check our computations.. This result will be tested against the 22nd PN order computation available in literature Fujita:2012cm , and results based on MST and numerical methods from the Black Hole Perturbation Toolkit (BHTK) BHPToolkit . The analytic expressions of all ingredients contributing to the 30PN expansion of the wave form and energy flux are collected in ancillary mathematical files. Numerical studies of unbounded systems are postponed to future investigations.
@@
This is the plan of the paper: in Section 2 we derive the PM and PN expansion of the confluent Heun function. In Section 3 we apply the results of Section 2 to the study of the gravitational waveforms emitted by a particle moving in a Schwarzschild background. In Section 4 we test our results against those obtained via the MST method and numerical integration. In Appendix A we derive the PM and PN expansions of the non-confluent Heun function. Section 5 contains our conclusions.
@@
By confluent Heun equation we will refer to the ordinary differential equation (6) of second order, with two regular singularities at $z=1,\infty$ and an irregular one at $z=0$. The equation is specified by five complex parameters dubbed here as $\{m_{1},m_{2},m_{3},u,x\}$. The expansion of the general solution of this equation in the ”soft limit” $x\ll 1$ is derived in section 2. These results apply to any physical problem governed by an equation of confluent Heun type. The focus of this paper is on the study of gravitational wave forms in Schwarzschild geometry. The Heun gravity dictionary in this case reads
@@ 2 The confluent Heun equation
${\rm Hc}^{\prime\prime}(t)+\left(\frac{\beta_{0}}{t}+\frac{\beta_{1}}{t-1}+\lambda\right){\rm Hc}^{\prime}(t)+\frac{\alpha_{1}t-\alpha_{0}}{t(t-1)}{\rm Hc}(t)=0$ (2)
@@
$Hc_{+}(t)$ $=$ ${\rm HeunC}[\alpha_{0},\alpha_{1},\beta_{0},\beta_{1},\lambda,t]$ (3)
$Hc_{-}(t)$ $=$ $t^{1-\beta_{0}}\text{HeunC}[\alpha_{0}+(1-\beta_{0})(\lambda-\beta_{1}),\alpha_{1}+\lambda(1-\beta_{0}),2-\beta_{0},\beta_{1},\lambda,t]$
@@ 2.1 Confluent Heun-Gauge dictionary
with $\delta=m_{1}m_{2}{+}m_{1}m_{3}{+}m_{2}m_{3}{-}m_{1}{-}m_{2}{-}m_{3}$, $x$ the gauge coupling and $u$ the Coulomb branch parameter. (6) is related to the standard form (2) of the Heun equation by the Heun gauge dictionary
$G_{1}(z)$ $=$ $z^{-{\alpha_{1}\over\lambda}}{\rm Hc}({\textstyle\frac{1}{z}})\qquad,\qquad t=z^{-1}\qquad,\qquad\lambda=-x$
@@
Equivalently, (6) can be written as a differential equation in the variable $y=x/z$, taking the form
@@
$G_{1}({\textstyle\frac{x}{y}})=g(x)G_{0}(y)$ (9)
where $g(x)$ is a $y$-independent function. In these variables, the confluent Heun equations (CHEs) (6) and (8) have irregular singularities at $z=0$ or $y=0$ and their solutions transform non-trivially under rotations around these points. We denote by $G^{0}_{\pm}(y)$, $G^{1}_{\pm}(z)$, the eigenvectors of this action, i.e. solutions of the CHEs satisfying
$G^{0}_{\alpha}(y\,e^{2\pi{\rm i}})=e^{-2\pi{\rm i}\alpha a}G^{0}_{\alpha}(y)\qquad,\qquad G^{1}_{\alpha}(z\,e^{2\pi{\rm i}})=e^{2\pi{\rm i}\alpha a}G^{1}_{\alpha}(z)$ (10)
with $\alpha=\pm$ and $a(u,x,m_{i})$ is an unknown function of the Heun parameters. The quantity $a$ will play a fundamental role in the following. One of the lessons of the Heun gravity correspondence is that the solutions of the CHE and the connection matrices look simpler if expressed in terms of $a$ rather than in terms of $u$.
We will focus on the "soft limit" where $x$ is taken small (large wave lengths), keeping finite either the variable $z$ or $y$. We will refer to these two limits as PM and PN respectively. The two choices lead to equations of hypergeometric type. Indeed, sending $x\to 0$ in (6) and (8) lead to hypergeometric equations with solutions
@@
where
@@
At this order, $a\underset{x\to 0}{\approx}\sqrt{u}$. For $x\neq 0$ but small one finds
${\rm PM}_{k}:$ $\qquad G^{0}_{\alpha}(y)=y^{\frac{1}{2}-\alpha a-m_{3}}\sum_{i=0}^{k}\sum_{j=0}^{\infty}c^{PM}_{ij}(\alpha a)x^{i}y^{j-i}+O(x^{k+1})$
${\rm PN}_{k}:$ $\qquad G^{1}_{\alpha}(z)=z^{-\frac{1}{2}+\alpha a+m_{3}}\sum_{j=0}^{k}\sum_{i=0}^{\infty}c^{PN}_{ij}(\alpha a)z^{i-j}x^{j}+O(x^{k+1})$ (13)
and
@@
with the coefficients $c^{PM}_{ij}(a)$, $c^{PN}_{ij}(a)$, $u_{i}(a)$ determined by solving the CHE equations order by order in $x$. Inverting (14) one finds the dependence of the monodromy eigenvalue $a(u)$ on the parameter $u$. Moreover, since $G^{1}_{\alpha}(z)$ and $G^{0}_{\alpha}({\textstyle\frac{x}{z}})$ are solutions of the same equation and have the same monodromy around $z=0$ they should be proportional to each other. We conclude therefore that in the overlap region, the so called near zone, where $x\ll z\ll 1$ we find
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.
@@ 2.2 PM approximation: exterior region
It is easy to check that $c_{ij}$, $\widehat{c}_{ij}$ are invariant under $a\to-a$. This is true at any PM order, so all the dependence on $\alpha$ of the solution is codified in $H^{0}_{\alpha}$. Explicit computations show that the coefficients $c_{i,i-1}$, $\widehat{c}_{i,i-1}$ can be always written in terms of derivatives of $u_{i}$. Indeed, one finds
@@
$\gamma_{m}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}=e^{-\partial_{m_{3}}{\cal F}_{\rm inst}}\quad$ (21)
with
${\cal F}_{\rm inst}\equiv-\sum_{i=1}^{\infty}{u_{i}\over i}x^{i}$ (22)